there is a ratio of 3 boys to every 4 girls. if there are 245 total preschoolers enrolled then how many of them are boys?
Answer by greenestamps(13215)



Given the ratio of 3 boys to every 4 girls,...

Let 3x = # of boys
Let 4x = # of girls

That is nearly always the best way to start on a problem like this where a ratio is given.

The total number of preschoolers is 245:

3x+4x=245
7x=245
x=245/7=35

ANSWER: The number of boys is 3x = 3(35) = 105